Technical Staff Blog

Archives June 2016

Last update on .

  •     New VPN configuration files are now available at
  •     Old VPN configuration files will expire on July 10th.
  •     A short VPN outage is scheduled for 12pm on June 24th.

openvpn.png We were unable to renew the VPN server certificate on schedule, and we would like to apologize for any inconvenience that this may ...

Last update on .

CIS Networking group is planning to make changes to the 10G switches, cit-cs-r14-sw-1 and cit-cs-r14-sw-2.  The switch config change requires a reboot.  They will be changing and rebooting the cit-cs-r14-sw-1 on Tuesday, June 21st at 5:30am and cit-cs-r14-sw-2 on Thursday, June 23rd at 6:00am.  The GPFS NSD nodes are connected to these switches ...

Last update on .

  • New VPN configuration files are now available at
  • Old VPN configuration files will expire on July 10th.
  • A short VPN outage is scheduled for 12pm on June 22nd.

 Tstaff will be renewing the Certificates used to provide the SSL connection for our VPN service.  As such, the new .ovpn configuration files ...

Last update on .

finger, floor, anyone, and the demise of rwho

Short version: The commands anyone and floor no longer work. finger/cslookup partly works but will no longer exist after this month. Its remaining functionality has been assumed by elbow.

Long version:

A week or so prior to the Great Precarious Filesystem Shift last month, we shut down our rwho service. Rwho was the system ...

Last update on .

EMC_Isilon.png On Saturday June 4th, CIS will be physically moving their EMC Isilon, the hardware that provides file services for the campus. The move is scheduled to commence at 6am and complete around 4pm. We currently have three production services relying on the Isilon: email lists, ownCloud, and time machine backups. In order to keep our ...

Last update on .

network-2400px.png This was supposed to be a trouble free morning... just power up the switch we moved downstairs yesterday and add our redundancy back in. It seems to be a trend with this project, things not going quite as we planned...

07:14: The second distribution switch was powered on at 7am. Unfortuantely, something happened that ...

Daily archives

Previous month

May 2016

Next month

July 2016